40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9199     Entry time: Thu Oct 3 22:14:06 2013
Author: Masayuki 
Type: Update 
Category: Green Locking 
Subject: Both arms ALS stabilized for IR resonance 

[Manasa, Masayuki]

We succeeded in stabilizing both the arms using ALS and get IR to resonate at the sametime.

Measurement

At each step we measured the _PHASE_OUT_Hz calibrated error signals for  Y in this configuration so as to get the in-loop noise of ALS control of YARM

1. we stabilized YARM off IR resonance by using ALS, misaligned ETMX, closed XARM green shutter. That means no IR flashing and no green in XARM.

2. we aligned the ETMX with XARM green shutter closed.

3. we opened the green shutter and locked the green laser with PDH to the XARM.

4. we stabilized the XARM using ALS and off resonance for IR.

5.We brought the XARM to IR resonance with YARM stabilized off IR resonance.

6. we brought the YARM to IR resonance

Beat frequencies when both the arms were stabilized and had IR resonating :
X arm beat frequency = 73.2 MHz; Y arm beat frequency = 26.6 MHz.

Attachment

1.the ALS in-loop noise in X and Y arms with IR off resonance and resonating.

2.the ALS in-loop noise in Y arm in each step from 1 to 6.(will follow soon)

Discussion

The Y arm ALS in-loop noise doesn't seem to be different in any of the configurations in step 1 to 6. This seems to mean that the ALS of the two arms are decoupled.

Actually we are not sure what changed from the last few days (when we were seeing some sort of coupling between the ALS of X and Y arm) except for YARM green PDH servo gain changed (see this entry),

Attachment 1: ALS_XY_inloop.pdf  86 kB  Uploaded Thu Oct 3 23:16:31 2013  | Hide | Hide all
ALS_XY_inloop.pdf
ELOG V3.1.3-